# Demystifying Platform Requirements for Diverse LLM Inference Use Cases

### v1

@@ VIII Extreme Scale Requirements: AI assistant
We envision an enormous 10T parameter Super-LLM model as the futuristic model from Table I. Our aim is to employ the Super-LLM as a standard AI assistant capable of engaging in real-time conversations with humans, furnishing answers on diverse topics, and retaining recollections of past interactions. Hence, supporting large context windows is imperative for AI assistants. This workload can be characterized as $S_{b}=4,\tau_{p}=\text{Variable},\tau_{d}=2000$. Given its role as a personal assistant, we would operate it with a batch size of 1. For seamless conversation, we assume it can generate output at the rate at which humans read or listen, which amounts to 300 words per minute[51].
(a) Memory Bandwidth (TB/s)
(b) Memory Capacity (PB).
@@
IX Related Works
LLM inference serving and analysis. We are seeing a massive uprising works related to LLM inference serving. There are works on optimizing the batching[130, 76, 43], memory optimization[61, 95] and scheduling[119, 67]. Various articles[18, 1, 6, 38, 7, 13, 19] provides metrics, analysis, insights, and best practices for LLM inference performance on current hardware systems. Various works provided a survey [131, 90, 58] of transformer inference and various optimizations in existing transformer architecture on current hardware systems.
HW for LLM inference and Hardware Modeling. Chiplet Cloud [109] is a chiplet-based ASIC AI supercomputer architecture that optimizes total cost of ownership (TCO) per generated token for serving large generative language models. LLMCompass[132] does DSE for LLM is its framework models the details of NPU architecture. There has been a lot of work in DNN inference design-space modeling[62, 117, 107, 94, 63, 98, 101, 113] which focus on modeling latency, bandwidth, and energy efficiency of various dataflows and loop ordering. Apart from compute modeling, network simulators like ASTRA-sim [111, 126] model communication in distributed training.
X Conclusion
With the recent advancement of LLMs, designing HW platform for efficient LLM serving becomes crucial to provide high throughput with low latency while maintaining cost-efficiency. To understand the implication of HW platform requirements for LLM inference considering diverse models, use cases, and SLO, we introduce GenZ, a first-order analytical tool. Using GenZ, we conduct three main analyses: 1) An analysis of the performance impact of different workload optimization, 2) An analysis of the performance impact of different platform characterizations, and 3) An analysis of the platform requirements for the target LLM inference with the target use case and SLO. We believe our proposed tool and analyses would shed light on future platform designs for next-generation HWs and LLMs.

### v2

@@ VII-E Extreme Scale Requirements: AI assistant
We envision an enormous 10T parameter Super-LLM model as the futuristic model from Table IV. Our aim is to employ the Super-LLM as a standard AI assistant capable of engaging in real-time conversations with humans, furnishing answers on diverse topics, and retaining recollections of past interactions. Hence, supporting large context windows is imperative for AI assistants. This workload can be characterized as $S_{b}=4,\tau_{p}=\text{Variable},\tau_{d}=2000$. Given its role as a personal assistant, we would operate it with a batch size of 1. For seamless conversation, we assume it can generate output at the rate at which humans read or listen, which amounts to 300 words per minute[110].
@@
VIII Related Works
LLM inference serving and analysis: We are seeing a massive uprising of works related to LLM inference serving. There are works on optimizing the batching [36, 34, 35], memory optimization [33, 70] and scheduling [111, 29]. Various articles [112, 113, 114, 115, 116, 117, 118] provides metrics, analysis, insights, and best practices for LLM inference performance on current hardware systems. Various works provided a survey  [13, 119, 120, 121, 122] of transformer inference and various optimizations in existing transformer architecture on current hardware systems.
Tools for AI Platform Modeling: Table I highlights key differences between GenZ and other performance models for LLMs. Simulators like ASTRA-sim [12], MadMax  [123] and vTrain  [124] focus on communication optimizations for distributed training and do not model different LLM architectures or LLM model optimizations on a distributed inference. Tools like LLM-viewer [13] provide ideal roofline estimation. Vidur [15] is an inference system simulator that focuses on studying the impact of different scheduling techniques on current hardware. LLMServingSim [16] provides a framework to compare scheduling algorithms. LLM-Compass [14] does DSE to optimize and generate ASIC configuration(Systolic arrays, cache sizes) for running dense models. There is a lack of a single tool/framework in the community that can help us study different LLM model architectures combined with the latest optimizations on distributed NPU platforms.
IX Extensions and Future Directions
Our work intentionally abstracts low-level microarchitectural details by encapsulating them as efficiency factors and using external tools for refined performance estimates. We also provide linear energy estimation and aggregated batching support, future works can extend these capabilities with advanced modelling (e.g. additional scheduling modules to simulate disaggregated /heterogeneous serving).
X Conclusions
We introduce GenZ, a framework with an indispensable capability of navigating the intricate design space of LLM inference, quantifying the interplay between various model and system-level optimizations, and steering the development of future AI platforms. We also show four key studies to present a roadmap for improving next-generation AI platform characteristics and selecting the appropriate architectural paradigms for running LLM inference. The source code is available at GitHub. GenZ can also be tried out on its website without any setup on your web browser.
We also provide linear energy estimation and aggregated batching support; future works can extend these capabilities with advanced modeling (e.g., additional scheduling modules to simulate disaggregated/heterogeneous serving).
XI Acknowledgment
This work was supported in part by CoCoSys, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. We thank the anonymous reviewers for their insightful feedback, which significantly improved the quality of this work. We are especially grateful to Amir Yazdanbakhsh for his valuable discussions and feedback, which helped shape the foundation of this research. We also thank Arijit Raychoudhury for his thoughtful discussions that were instrumental in realizing the full potential of this work.